# CBMC-V3: A CNS-inspired Control Framework Towards Manipulation Agility with SNN

### v1

@@ Abstract
As robotic arm applications extend beyond industrial settings into healthcare, service, and daily life, existing control algorithms struggle to achieve the agile manipulation required for complex environments with dynamic trajectories, unpredictable interactions, and diverse objects. This paper presents a biomimetic control framework based on Spiking Neural Networks (SNN), inspired by the human Central Nervous System (CNS), to achieve agile control in such environments. The proposed framework features five control modules (cerebral cortex, cerebellum, thalamus, brainstem, spinal cord), three hierarchical control levels (first-order, second-order and third-order), and two information pathways (ascending and descending). Each module is fully implemented using SNN. The spinal cord module uses spike encoding and Leaky Integrate-and-Fire (LIF) neurons for feedback control. The brainstem module employs a network of LIF and non-spiking LIF neurons to dynamically adjust spinal cord parameters via reinforcement learning. The thalamus module similarly adjusts the cerebellum’s torque outputs. The cerebellum module uses a recurrent SNN to learn the robotic arm’s dynamics through regression, providing feedforward gravity compensation torques. The framework is validated both in simulation and on real-world robotic arm platform under various loads and trajectories. Results demonstrate that our method outperforms the industrial-grade position control in manipulation agility.
@@ I Introduction
With the advancement of actuator and material technologies, robotic arms have gradually increased in degrees of freedom while decreasing in cost, enabling them to perform more delicate operations. Consequently, their applications have expanded in recent years to fields such as healthcare, services, and entertainment. As robotic arms become increasingly integrated into our daily lives, the complexity of their operating environments and the agility required for diverse, fine-grained tasks pose new challenges to control strategies. For instance, a 7-DOF robotic arm can readily cut a circular contour on a workpiece or transport heavy objects, yet it still struggles to perform a task as subtle as peeling an apple—a task easily accomplished by the human arm. Compared with biological systems, even the most advanced robotic arms today fall short of the agility of the human upper limb [mason2018], underscoring the existence of advanced motor control mechanisms in humans that remain underutilized in robotic systems.
@@
In well-modeled environments, control algorithms combining Whole-Body Control (WBC) [Abe2007, liu2011interactive, weighttans2011, Lee2012Intermediate, Escande2014, Kim2019, han2021recursive] and Model Predictive Control (MPC) [carron2019data, pankert2020perceptive, wang2023trajectory] can effectively counteract disturbances and achieve strong control performance. Such model-based approaches reflect human capabilities in applying prior knowledge to solve specific tasks. However, they lack the adaptability inherent to human motor learning. In complex scenarios, variations in the size, mass, or dynamics of manipulated objects may render the kinematic and dynamic models inaccurate, limiting the effectiveness of purely model-based strategies.
@@
With the rapid progress of neural networks, their powerful approximation and learning capabilities have motivated researchers to employ them for adaptive model correction in response to environmental changes. For example, [chaoui2009ann] designed feedforward and feedback neural networks to compensate for joint flexibility and friction; [he2017adaptive] employed fuzzy neural networks with adaptive impedance control to handle unknown dynamics and state constraints; [wang2018adaptive] proposed a neural-network-based adaptive controller capable of addressing nonlinear dynamics and external disturbances without prior models; [salloom2020adaptive] integrated genetic algorithm observers with neural networks to compensate for underwater disturbances; [liu2019adaptive] investigated the role of hidden-layer size in adaptive control and proposed dynamic adjustment strategies; and [pham2020adaptive] applied radial basis function networks with dynamic surface control for robust dual-arm coordination. While these approaches achieved progress, they rely on artificial neural networks (ANN), which inherently disregard the temporal-spatial characteristics of motion, focusing only on instantaneous control. Furthermore, they remain constrained by robotic dynamics models, with network structures and training methods often tailored to specific problems and tightly coupled to complex mathematical formulations, thereby limiting generalizability.
@@
Recently, spiking neural networks (SNN) combined with brain-inspired architectures have emerged as promising alternatives. By emulating biological neurons, SNNs process and store information through spike trains and membrane potential dynamics, offering advantages in computational speed, energy efficiency, information dimension, and biological plausibility, making them well-suited for robotic control [bing2018survey]. For instance, [carrillo2008real] developed a cerebellum-inspired SNN trained via spike-timing-dependent plasticity (STDP) to learn manipulator dynamics for torque control; [abadia2019robot] extended this model to human–robot collaboration scenarios, demonstrating robustness under unknown disturbances; [abadia2021cerebellar] further established resilience to control delays, highlighting potential in teleoperation and cloud robotics and refines this framework to enable robotic arms to dynamicallyy adjust compliance. However, these methods suffer from high structural complexity, computational cost, hyperparameter sensitivity, and limited interpretability, as they focus narrowly on cerebellar functionality while neglecting the integrative role of the broader central nervous system (CNS), which coordinates control through interactions among several neural areas[Halassa_2022].
@@ II-A Module function
The brainstem has been shown to contain pathways involved in movement termination and reorientation, and to interact with the spinal cord to coordinate motor behaviors [merel2019hierarchical, annurev:/content/journals/10.1146/annurev-neuro-082321-025137]. Motivated by this, the brainstem module in our framework dynamically adjusts the PD gains for each joint based on trajectory information,i.e.,
@@
providing the spinal cord module with time-varying proportional and derivative coefficients, i.e., $k_{p,i}$ and $k_{v,i}$ for $i=1,\dots,n$. By integrating spinal feedback control with brainstem-driven adaptive gain modulation, the framework circumvents the complexity of manipulator dynamics while compensating for structured uncertainties through learning.
@@
where $\bm{\tau}_{\text{fb}}(\bm{\dot{q}}_{d},\bm{q}_{d},\bm{\dot{q}},\bm{q})$ is produced by the spinal cord and brainstem module, $\bm{\tau}g(\bm{q},\bm{F}_{\text{ef}})$ is produced by the cerebellum and thalamus module, and $(\tau_{1},\dots,\tau_{n})^{T}$ represents the joint torques delivered to the actuators.
@@ II-B Hierarchical control level
  • Second-order level: Including the brainstem and thalamus module, this level operates at an intermediate frequency. It serves as a relay between the first- and third-order levels, performing two main functions: regulating the feedback control coefficients of the first-order loop and integrating and filtering the torques provided by the third-order layer to supply feedforward torque to the first-order loop.
@@ II-D A complete operational cycle
In each cycle, proprioceptive data (actual joint states $(\bm{\dot{q}},\bm{q})$) and desired joint states $(\bm{\dot{q}}_{d},\bm{q}_{d})$ are transmitted via first-order neurons to the spinal cord. Part of this input is locally processed to generate feedback torque $\bm{\tau}_{\text{fb}}$; part is relayed via second-order neurons to the brainstem, which adaptively provides PD gains $(\bm{K}_{P},\bm{K}_{V})$; and part is sent via third-order neurons to the cerebellum, which outputs gravity compensation matrix $\bm{T}_{\text{cb}}$. These ascending transmissions complete the flow of proprioceptive information from sensors to the modules.
@@ III-A Spiking neuron
$i\left(t\right)=\int_{0}^{\infty}{S\left(s-t\right)exp\left(-s/\tau_{s}\right)ds}$ (7)
@@
$\tau_{m}\frac{du}{dt}=u_{\text{reset}}-u\left(t\right)+R\left(i_{0}\left(t\right)+\sum w_{j}i_{j}\left(t\right)\right)$ (8)
@@
In our framework, we adopt a backpropagation-based training method[8891809] for SNN. Since the activation function of the neuron in 10 is a discontinuous function, directly computing its derivative can lead to network instability. Consequently, various surrogate gradient methods have been proposed [8891809]. A common approach is to use the original step function $y_{i}=\Theta(a_{i})$ during the forward pass, while replacing its derivative with $\theta^{\prime}(a_{i})=\sigma^{\prime}(a_{i})$ instead of $\Theta^{\prime}(a_{i})$ during backpropagation, where $\sigma(a_{i})$ is referred to as the surrogate activation function. Typically, $\sigma(a_{i})$ is a smooth and continuous function that resembles the shape of $\Theta(a_{i})$ while enabling stable gradient computation.
@@ III-B Cerebellum module
Figure 2: (a) Structure of one network unit during training in the cerebellum module. (b) Structure of cerebellum network during prediction.
@@
After training, the network can operate without reference torques, with CF neurons no longer required, relying solely on the learned synaptic weights to generate gravity compensation torques. To accommodate different loads, two network units are trained separately under light (no load) and heavy (3 kg load) modes using the supervised learning procedure described in Algorithm 2. These trained networks are then combined to form the complete cerebellum module structure illustrated in Fig. 2b, enabling adaptive gravity compensation across multiple load conditions.
In summary, after training, the cerebellum module receives as input the actual joint positions of the robotic arm, $\bm{q}\in\mathbb{R}^{7}$, and outputs a gravity compensation matrix composed of two torque vectors, $\bm{\tau}_{\text{cb,L}}\in\mathbb{R}^{7}$ and $\bm{\tau}_{\text{cb,H}}\in\mathbb{R}^{7}$, i.e., $\bm{T}_{\text{cb}}=(\bm{\tau}_{\text{cb,L}},\bm{\tau}_{\text{cb,H}})$. This matrix corresponds to the learned patterns of the cerebellum module and provides gravity compensation for trajectory tracking tasks under both light and heavy conditions. The outputs are subsequently integrated by the thalamus module to achieve gravity compensation for arbitrary loads within a specified range. During trajectory tracking, the prediction process of the cerebellum module is described in Algorithm 3, which can be implemented as Sub-thread 4 in Algorithm 1.
@@ III-C Thalamus module
The thalamus module is implemented as a three-layer fully connected SNN, as illustrated in Fig. 3. Its input, $\bm{F}_{\text{ef}}\in\mathbb{R}^{6}$, corresponds to the six-dimensional force/torque sensor data measured at the end-effector, while its output, $\bm{w}=(w_{1},w_{2})^{T}\in\mathbb{R}^{2}$, represents the weight coefficients assigned to the two gravity compensation torques provided by the cerebellum module.
@@ III-E Spinal cord module
Following this encoding process, two sets of discrete spike trains are generated—one representing position information and the other representing velocity information. These spike trains are then transmitted to stimulate the activity of non-spiking LIF neurons in the output layer, with stimulation strength modulated by the feedback control gains $k_{p,i}$ and $k_{v,i}$ provided by the brainstem module, as illustrated by the blue and red dashed lines, respectively. Accordingly, at each time step, the membrane potential increments of the output non-spiking LIF neurons can be expressed as
@@ IV Simulation analysis
The algorithm is implemented in Python (version 3.12.4), and all neural networks are constructed using PyTorch [paszke2019pytorch] (version 2.4.0). The spiking neurons involved in the algorithm are implemented with the SNN learning framework SpikingJelly [doi:10.1126/sciadv.adi1480] (version 0.0.0.0.14). The simulation environment employed is PyBullet [coumans2021] (version 3.2.6), an open-source robotics simulation tool whose underlying physics engine is based on the open-source Bullet Physics SDK. The controlled robotic arm consists of seven actuated joints, seven links and a six-dimensional torque sensor mounted at the flange of the end-effector. The simulation time step is set to 1 ms, and the control frequency is configured to 1 kHz.
@@
$\text{RMSE}=\frac{1}{N}\sum_{t=1}^{N}RMSE(t)$ (27)
@@ IV-C Framework validation
To verify this correspondence, ablation studies were conducted by selectively disabling individual modules, as shown in Fig. 6e, 6f and 6g. Removing the brainstem’s learning capability by fixing feedback control coefficients reduces the number of peaks to two and delays the first peak, indicating diminished adaptability to load change. Similarly, disabling the thalamus module by fixing mode-combination weights eliminates its characteristic peak, leaving only brainstem and cerebellar responses. Finally, replacing the cerebellum’s adaptive torque generation with fixed compensation suppresses the cerebellar peak and weakens the thalamic one, since the thalamus relies on cerebellar outputs to optimize weight assignment. These results confirm the functional contributions of each module and their hierarchical cooperation.
In summary, the proposed framework exhibits robust adaptability and rapid convergence to load variations, with distinct loop-specific responses emerging in a biologically consistent sequence. This layered activation mirrors the progression of human neural responses to perturbations, namely M1 reflexes, M2 reflexes, and voluntary movements [forgaard2015voluntary], thereby reinforcing the plausibility and effectiveness of the neuromorphic design.
@@ V Experiment results
$RMSE_{p}=\frac{1}{N}\sum_{t=1}^{N}RMSE_{p}(t)$ (31)
$RMSE_{o}=\frac{1}{N}\sum_{t=1}^{N}RMSE_{o}(t)$ (32)

### v2

@@ Abstract
As robotic arm applications extend beyond industrial settings into service-oriented sectors such as catering, household and retail, existing control algorithms struggle to achieve the agile manipulation required for complex environments with dynamic trajectories, unpredictable interactions, and diverse objects. This paper presents a biomimetic control framework based on Spiking Neural Networks (SNN), inspired by the human Central Nervous System (CNS), to achieve agile control in such environments. The proposed framework features five control modules (cerebral cortex, cerebellum, thalamus, brainstem, and spinal cord), three hierarchical control levels (first-order, second-order, and third-order), and two information pathways (ascending and descending). Each module is fully implemented using SNN. The spinal cord module uses spike encoding and Leaky Integrate-and-Fire (LIF) neurons for feedback control. The brainstem module employs a network of LIF and non-spiking LIF neurons to dynamically adjust spinal cord parameters via reinforcement learning. The thalamus module similarly employs a network of LIF and non-spiking LIF neurons to adjusts the cerebellum’s torque outputs via reinforcement learning. The cerebellum module, which provides feedforward gravity compensation torques, uses a recurrent SNN to learn the robotic arm’s dynamics through regression. The framework is validated both in simulation and on real-world robotic arm platform under various loads and trajectories. Results demonstrate that our method outperforms the industrial-grade position control in manipulation agility.
@@ I Introduction
In recent years, robotic arms are increasingly applied in service-oriented sectors such as catering, household and retail. Compared with traditional industrial settings, these scenarios feature more complex environments and diverse tasks while requiring better agility, thereby posing new challenges to control strategies. For instance, a 7-DOF robotic arm can readily cut a circular contour on a workpiece or transport heavy replacement parts, yet it still struggles to perform a task as subtle as peeling an apple, which can be easily accomplished by the human arm. Compared with biological systems, even the most advanced robotic arms today fall short of the agility of human upper limb [mason2018], underscoring the existence of advanced motor control mechanisms in humans that remain underutilized in robotic systems.
@@
In well-modeled environments, control algorithms combining Whole-Body Control (WBC) [Abe2007, liu2011interactive, weighttans2011, Lee2012Intermediate, Escande2014, Kim2019, han2021recursive] and Model Predictive Control (MPC) [carron2019data, pankert2020perceptive, wang2023trajectory] can effectively counteract disturbances and achieve strong control performance. Such model-based approaches reflect human capabilities in applying prior knowledge to solve specific tasks. However, they lack the adaptability inherent to human motor learning. In complex scenarios, variations in the size, mass, or materials of manipulated objects may render the kinematic and dynamic models inaccurate, limiting the effectiveness of these model-based strategies.
@@
With the rapid progress of neural networks, their powerful approximation and learning capabilities have motivated researchers to employ them for adaptive model correction in response to environmental changes. For example, [chaoui2009ann] designed feedforward and feedback neural networks to compensate for joint flexibility and friction; [he2017adaptive] employed fuzzy neural networks with adaptive impedance control to handle unknown dynamics and state constraints; [wang2018adaptive] proposed a neural-network-based adaptive controller capable of addressing nonlinear dynamics and external disturbances without prior models; [salloom2020adaptive] integrated genetic algorithm observers with neural networks to compensate for underwater disturbances; [liu2019adaptive] investigated the role of hidden-layer size in adaptive control and proposed dynamic adjustment strategies; and [pham2020adaptive] applied radial basis function networks with dynamic surface control for robust dual-arm coordination. Despite these progression, they rely on artificial neural networks (ANN), which inherently disregard the temporal-spatial characteristics of motion, focusing only on instantaneous control. Furthermore, they remain constrained by robotic dynamics models, with network structures and training methods often tailored to specific problems and tightly coupled to complex mathematical formulations, thereby limiting generalizability.
@@
Recently, spiking neural networks (SNN) combined with brain-inspired architectures have emerged as promising alternatives. By emulating biological neurons, SNNs process and store information through spike trains and membrane potential dynamics, offering advantages in computational speed, energy efficiency, information dimension, and biological plausibility, making them well-suited for robotic control [bing2018survey]. For instance, [carrillo2008real] developed a cerebellum-inspired SNN trained via spike-timing-dependent plasticity (STDP) to learn manipulator dynamics for torque control; [abadia2019robot] extended this model to human–robot collaboration scenarios, demonstrating robustness under unknown disturbances; [abadia2021cerebellar] further established resilience to control delays, highlighting potential in teleoperation and cloud robotics; [doi:10.1126/scirobotics.adp2356] refines this framework to enable robotic arms to dynamically adjust compliance. However, these methods suffer from structural complexity, computational cost, hyperparameter sensitivity, and limited interpretability, as they focus narrowly on cerebellar functionality while neglecting the integrative role of the broader central nervous system (CNS), which coordinates control through interactions among several neural areas[Halassa_2022].
@@ II-A Module function
The brainstem has been shown to contain pathways involved in movement termination and reorientation, and to interact with the spinal cord to coordinate motor behaviors [merel2019hierarchical, annurev:/content/journals/10.1146/annurev-neuro-082321-025137]. Motivated by this, the brainstem module in our framework dynamically adjusts the PD gains for each joint based on trajectory information, providing the spinal cord module with time-varying proportional and derivative coefficients, i.e.,
@@
By integrating spinal feedback control with brainstem-driven adaptive gain modulation, the framework circumvents the complexity of manipulator dynamics while compensating for structured uncertainties through learning.
@@
where $\bm{\tau}_{\text{fb}}(\bm{\dot{q}}_{d},\bm{q}_{d},\bm{\dot{q}},\bm{q})$ is produced by the spinal cord and brainstem module, $\bm{\tau}_{g}(\bm{q},\bm{F}_{\text{ef}})$ is produced by the cerebellum and thalamus module, and $(\tau_{1},\dots,\tau_{n})^{T}$ represents the joint torques delivered to the actuators.
@@ II-B Hierarchical control level
  • Second-order level: Including the brainstem and thalamus module, this level operates at an intermediate frequency. It serves as a relay between the first-order and third-order levels, performing two main functions: regulating the feedback control coefficients of the first-order level and integrating and filtering the torques provided by the third-order level to supply feedforward torque to the first-order level.
@@ II-D A complete operational cycle
In each cycle, proprioceptive data (actual joint states $(\bm{\dot{q}},\bm{q})$ and desired joint states $(\bm{\dot{q}}_{d},\bm{q}_{d})$) are transmitted via first-order neurons to the spinal cord. Part of this input is locally processed to generate feedback torque $\bm{\tau}_{\text{fb}}$; part is relayed via second-order neurons to the brainstem, which adaptively provides PD gains $(\bm{K}_{P},\bm{K}_{V})$; and part is sent via third-order neurons to the cerebellum, which outputs gravity compensation matrix $\bm{T}_{\text{cb}}$. These ascending transmissions complete the flow of proprioceptive information from sensors to the modules.
@@ III-A Spiking neuron
$i\left(t\right)=\int_{0}^{\infty}{S\left(s-t\right)\exp\left(-s/\tau_{s}\right)\text{d}s}$ (7)
@@
$\tau_{m}\frac{\text{d}u}{\text{d}t}=u_{\text{reset}}-u\left(t\right)+R\left(i_{0}\left(t\right)+\sum w_{j}i_{j}\left(t\right)\right)$ (8)
@@
In our framework, we adopt a backpropagation-based training method[8891809] for SNN. Since the activation function of the neuron in equation (10) is a discontinuous function, directly computing its derivative can lead to network instability. Consequently, various surrogate gradient methods have been proposed [8891809]. A common approach is to use the original step function $y=\Theta(a)$ during the forward pass, while replacing its derivative with $\theta^{\prime}(a)=\sigma^{\prime}(a)$ instead of $\Theta^{\prime}(a)$ during backpropagation, where $\sigma(a)$ is referred to as the surrogate activation function. Typically, $\sigma(a)$ is a smooth and continuous function that resembles the shape of $\Theta(a)$ while enabling stable gradient computation.
@@ III-B Cerebellum module
Figure 2: (a) Structure of one network unit in the cerebellum module during training. (b) Structure of cerebellum network during prediction.
@@
After training, the network can operate without reference torques, with CF neurons no longer required, relying solely on the learned synaptic weights to generate gravity compensation torques. To accommodate different loads, two network units are trained separately under light (no load) and heavy (3 kg load) modes using the supervised learning procedure described in Algorithm 2. These trained networks are then combined to form the complete cerebellum module structure illustrated in Fig. 2b.
In summary, after training, the cerebellum module receives as input the actual joint positions of the robotic arm, $\bm{q}\in\mathbb{R}^{7}$, and outputs a gravity compensation matrix composed of two torque vectors, i.e., $\bm{T}_{\text{cb}}=(\bm{\tau}_{\text{cb,L}},\bm{\tau}_{\text{cb,H}})$, $\bm{\tau}_{\text{cb,L}}\in\mathbb{R}^{7}$ and $\bm{\tau}_{\text{cb,H}}\in\mathbb{R}^{7}$. This matrix corresponds to the learned patterns of the cerebellum module and provides gravity compensation for trajectory tracking tasks under both light and heavy conditions. The outputs are subsequently integrated by the thalamus module to achieve gravity compensation for arbitrary loads within a specified range. During trajectory tracking, the prediction process of the cerebellum module is described in Algorithm 3, which can be implemented as Sub-thread 4 in Algorithm 1.
@@ III-C Thalamus module
The thalamus module contains five layers of neurons, as illustrated in Fig. 3. Its input $\bm{F}_{\text{ef}}\in\mathbb{R}^{6}$ and $\bm{T}_{\text{cb}}=(\bm{\tau}_{\text{cb,L}},\bm{\tau}_{\text{cb,H}})\in\mathbb{R}^{7\times 2}$, corresponds to the six-dimensional force/torque sensor data measured at the end-effector, and the cerebellum module output, respectively. Its output, $\bm{w}=(w_{1},w_{2})^{T}\in\mathbb{R}^{2}$, represents the weight coefficients assigned to the two gravity compensation torques provided by the cerebellum module.
@@ III-E Spinal cord module
Following this encoding process, two sets of discrete spike trains are generated—one representing position information and the other representing velocity information. These spike trains are then transmitted to stimulate the activity of non-spiking LIF neurons in the output layer, with stimulation strength modulated by the feedback control gains $k_{p,i}$ and $k_{v,i}$, $i=1,\cdots,7$ provided by the brainstem module, as illustrated by the blue and red dashed lines, respectively. Accordingly, at each time step, the membrane potential increments of the output non-spiking LIF neurons can be expressed as
@@ IV Simulation analysis
The algorithm is implemented in Python (version 3.12.4), and all neural networks are constructed using PyTorch [paszke2019pytorch] (version 2.4.0). The spiking neurons involved in the algorithm are implemented with the SNN learning framework SpikingJelly [doi:10.1126/sciadv.adi1480] (version 0.0.0.0.14). The simulation environment employed is PyBullet [coumans2021] (version 3.2.6), an open-source robotics simulation tool whose underlying physics engine is based on the open-source Bullet Physics SDK. The controlled robotic arm consists of seven actuated joints, seven links and a six-dimensional torque sensor mounted at the flange of the end-effector. The control cycle of each layer is set as $T_{1}=1$ms, $T_{2}=10$ms, $T_{3}=50$ms. The simulation time step is set to 1 ms, and the control frequency is configured to 1 kHz.
@@
$\text{RMSE}=\frac{1}{N}\sum_{t=1}^{N}\text{RMSE}(t)$ (27)
@@ IV-C Framework validation
To verify this correspondence, ablation experiments were conducted by selectively disabling individual modules, as shown in Fig. 6e, 6f and 6g. Removing the brainstem’s learning capability by fixing feedback control coefficients reduces the number of peaks to two and delays the first peak, indicating diminished adaptability to load change. Similarly, disabling the thalamus module by fixing mode-combination weights eliminates its characteristic peak, leaving only brainstem and cerebellar responses. Finally, replacing the cerebellum’s adaptive torque generation with fixed compensation suppresses the cerebellar peak and weakens the thalamic one, since the thalamus relies on cerebellar outputs to optimize weight assignment. These results confirm the functional contributions of each module and their hierarchical cooperation.
In summary, the proposed framework exhibits robust adaptability and rapid convergence to load variations, with distinct loop-specific responses emerging in a biologically consistent sequence. This step-by-step activation phenomenon mirrors the progression of human neural responses to perturbations, namely M1 reflexes, M2 reflexes, and voluntary movements [forgaard2015voluntary], thereby reinforcing the plausibility and effectiveness of the neuromorphic design.
@@ V Experiment results
$\text{RMSE}_{p}=\frac{1}{N}\sum_{t=1}^{N}\text{RMSE}_{p}(t)$ (31)
$\text{RMSE}_{o}=\frac{1}{N}\sum_{t=1}^{N}\text{RMSE}_{o}(t)$ (32)